OpenStudy (anonymous):

what would be the two equations for (-3,-6) and (1,6)??

OpenStudy (compassionate):

y = mx + b Plug them in. The first one, x = -3 and y = -6 -6 = -3m + b

OpenStudy (anonymous):

thank you so much, you actually show me how to do it.. i think i understand now..